OpenStudy (anonymous):

Mindy has a 20-year adjustable rate mortgage with a rate of 3.6% for the first 4 years. The monthly payment is $1,375.01. The amount of the mortgage is $235,000. What is the remaining balance after 4 years rounded to the nearest dollar? $188,000 $200,464 $169,000 $191,026

OpenStudy (tkhunny):

Can you use most if the information to recalculate the initial payment?

OpenStudy (anonymous):

Can I use 1,375.01 and 235,000

OpenStudy (tkhunny):

Nope. Use everything except the payment and recalculate the payment. 20 years 3.6% $235,000

OpenStudy (anonymous):

I plug it into this M=B[((i)(1+i)^(n*t))/((1+i)^(n*t)-(1))

OpenStudy (tkhunny):

That should be the right formula. See if you match my results. i = 0.036 j = I/12 v = 1/(1+j) n = 240 I use this one: \(235000 = Pmt\cdot\dfrac{v-v^{n+1}}{1-v}\) This creates Pmt = 1375.01 The solution is to work the formula a second time Without the Total Amount. This is what we need to find. With the Payment: 1375.01 With n = 16*12 = 192 (4 years less)

OpenStudy (anonymous):

would I=r/n ?

OpenStudy (tkhunny):

No, the definitions are provided above for my version. The version from your book or course materials should have similar versions. j = i/n = 0.036/12 in this case. Your formula doesn't use 'v', which, in my opinion, make life simpler.

OpenStudy (anonymous):

@tkhunny I got 1068.42

OpenStudy (amistre64):

loan amount: 235 000 interest to be compounded:3.6% monthly payments: 1 375.01 after 4 years (48 months) the balance remaing is just: Bk^n - P(1-k^n)/(1-k) 235000k^(48)-1375.01(1-k^(48))/(1-k), k=1+.036/12

OpenStudy (amistre64):

all they are asking you is what the remaining balance would be in order to restructure the loan after the initial period ends; or if this was a balloon mortgage ... the amount of the balloon payment